OpenStudy (anonymous):

What is the standard form of the equation of the given circle with center A(4, 8)?

OpenStudy (anonymous):

A) (x − 4)2 + (y − 8)2 = 2.5 B) (x + 4)2 + (y + 8)2 = 6.25 C) (x + 4)2 + (y + 8)2 = 2.5 D) (x − 4)2 + (y − 8)2 = 6.25

OpenStudy (anonymous):

(x-a)^2 + (y-b)^2 = r^2 hence its